# On Spheres with $k$ Points Inside

### v2

@@ 3.1 Worpitzky’s Identity for Eulerian Numbers
As an application of our covering results, we show how the multiplicities from Theorem 2.3 are related to the volumes of hypersimplices and to Eulerian numbers. Specifically, we show that de Laplace’s relation for hypersimplices [7] implies Worpitzky’s identity for Eulerian numbers [13], and vice versa.
We begin by introducing the three main concepts we need in this subsection. Letting $d$ be a positive integer, the number of descents in a permutation $j_{1},j_{2},\ldots,j_{d}$ of $1,2,\ldots,d$ is the number of indices $11\leq i\leq d-$ 1 such that $j_{i}>j_{i+1}$. For $10\leq k\leq d-$ 1, the Eulerian number for $d$ and $k$ is the number of permutations with exactly $k$ descents. For example, $A(d,0)=A(d,d-1)=1$, for every $d$, and $\sum_{k=0}^{d-1}A(d,k)=d!$. Less obvious is Worpitzky’s identity [13] between two polynomials from more than a century ago:
@@
Next, let $x_{0},x_{1},\ldots,x_{d}$ be $d+1$ affinely independent points and write $\Delta={{\rm conv\,}{\{x_{0},x_{1},\ldots,x_{d}\}}}$ for the $d$-simplex they span. For each $1\leq p\leq d$, the convex hull of the $p$-fold sums of the points is a $d$-dimensional convex polytope referred to as a $d$-hypersimplex of order $p$, denoted $\Delta_{d}^{(p)}$. Clearly, $\Delta_{d}^{(1)}=\Delta$, and more generally, $\Delta_{d}^{(p)}$ is a homothetic copy of the convex hull of the barycenters of all $(p-1)$-dimensional faces of $\Delta$. Since the barycenter of a $(p-1)$-simplex is $\nicefrac{{1}}{{p}}$ times the sum of its vertices, the volume of that polytope is $\nicefrac{{1}}{{p^{d}}}$ times the volume of the homothetic hypersimplex. Define the relative volume of $\Delta_{d}^{(p)}$ as $v(d,p)={\operatorname{vol}_{d}(\Delta^{(p)}_{d})}/{\operatorname{vol}_{d}(\Delta)}$, and observe that it does not depend on the choice of $\Delta$. Again more than a century ago, de Laplace [7] proved that these relative volumes are Eulerian numbers:
@@
see also the combinatorial proof of the same equation by Stanley [10]. The third concept is the dual of the order-$n$ Voronoi tessellation of a finite set $A\subseteq{{\mathbb{R}}}^{d}$, introduced in 1990 by Aurenhammer [2]. Referring to this dual as the order-$n$ Delaunay mosaic of $A$, denoted ${{\rm Del}_{n}{({A})}}$, it is defined by its $d$-cells, each the convex hull of a collection of averages of $n$ points selected from $A$. Specifically, for each $1\leq p\leq d$ and every $(n-p)$-hefty simplex, take all sets of cardinality $n$ that contain all $n-p$ points inside the circumsphere together with any $p$ points on the circumsphere. E.g. for $p=1$, we get $d+1$ averages whose convex hull is a homothetic copy of the original $d$-simplex, and for $p=2$, we get a homothetic copy of the convex hull of the midpoints of the edges of the $d$-simplex. Collecting these polytopes, we get ${{\rm Del}_{n}{({A})}}$. For $n=1$, we have ${{\rm Del}_{1}{({A})}}={{\rm Del}{({A})}}$, and more generally ${{\rm Del}_{n}{({A})}}$ has a $d$-cell for every $(n-p)$-hefty simplex in which $p$ varies from $1$ to $d$. Since the vertices are averages of $n$ points, each $d$-cell in ${{\rm Del}_{n}{({A})}}$ has volume $\nicefrac{{1}}{{n^{d}}}$ times the volume of the corresponding hypersimplex. It is now easy to prove the following relation for the relative volumes of the hypersimplices.
@@
in which $\binom{n+d-p}{n-p}=0$ whenever $n-p<0$.
Proof 3.2.
Let $A$ be any Delaunay set—and not just a thin Delaunay set—in ${{\mathbb{R}}}^{d}$, so that every ball whose radius exceeds some given constant contains at least one point of $A$. Let $R>0$ be sufficiently large and consider all $d$-cells in ${{\rm Del}_{n}{({A})}}$ that are contained in $[-R,R]^{d}$. Setting $n^{\prime}=\max\{0,n-d\}$, there are $n-n^{\prime}\leq d$ different types of $d$-cells to be considered, namely homothetic copies of hypersimplices of orders $1$ to $n-n^{\prime}$ defined by $(n-p)$-hefty simplices for $1\leq p\leq n-n^{\prime}$. The total volume of these $d$-cells is $(2R)^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$, since we miss only a constant width neighborhood of each facet of the hypercube.
Consider now an $(n-p)$-hefty simplex of $A$. Generically, its circumsphere passes through $d+1$ points and encloses $n-p$ of the points in $A$. By definition of Delaunay set, the radius of this circumsphere is bounded from above by a constant times $n-p$. Furthermore, the $(n-p)$-hefty simplex contains every $d$-cell in ${{\rm Del}_{n}{({A})}}$ it may determine since the vertices of the latter are averages of the vertices of the $(n-p)$-hefty simplex. By Theorem 2.3, the $(n-p)$-hefty simplices that define $d$-cells of ${{\rm Del}_{n}{({A})}}$ inside the hypercube therefore cover most of the hypercube exactly $\binom{n+d-p}{n-p}$ times. It follows that the total volume of these $p$-hefty simplices is $\binom{n+d-p}{n-p}(2R)^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$. By definition of relative volume, the total volume of the corresponding hypersimplices thus is ${v(d,p)}\binom{n+d-p}{n-p}(2R)^{d}/n^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$. Taking the sum for $1\leq p\leq n-n^{\prime}$, dividing by $(2R)^{d}$, and taking the limit as $R$ goes to infinity, we get the claimed relation.
To see that Theorem 3.1 together with de Laplace’s relation implies Worpitzky’s identity and together with Worpitzky’s identity implies de Laplace’s relation, we change the summation index in (4) from $p$ to $k=d-p$ and apply $\binom{n+k}{n-d+k}=\binom{n+k}{d}$ to get
@@
Since this relation holds for every positive integer, $n$, it also holds if we treat $\binom{n+k}{d}$ and $n^{d}$ as polynomials of degree $d$ in $n$. Substituting $A(d,k)=A(d,d-k-1)$ for $v(d,d-k)$ using de Laplace’s relation (3), we get Worpitzky’s identity (2). To see the other direction, we observe that the polynomials given by the binomial coefficients are linearly independent, so there is only one way to write $n^{d}$ as their linear combination, and it is given by Worpitzky’s identity. Comparing (5) with (2), we get $v(d,d-k)=A(d,k)=A(d,d-k-1)$, which is (3).
3.2 $k$-sets and $k$-facets
In this section, we briefly discuss connections between the covering numbers studied in Section 2 and the $k$-sets and $k$-facets studied in discrete geometry; see e.g. [8, Chapter 11]. Letting $A$ be a generic set of $n$ points in ${{\mathbb{R}}}^{d}$, a $k$-set is a set of $k$ points, $B\subseteq A$, such that $B$ and $A\setminus B$ can be separated by a hyperplane, and a $k$-facet is a set of $d$ points, $\Delta\subseteq A$, such that the hyperplane passing through these points partitions $A\setminus\Delta$ into $k$ and $n-d-k$ points on its two sides. We refer to [11, Section 2.2] for a discussion of the relation between these two concepts.
We present alternative proofs of two well-established results, which we state in terms of $k$-facets. Both proofs make use of the inversion of $A$ through the unit sphere centered at a point $x\in{{\mathbb{R}}}^{d}\setminus A$, which maps a point $a\in{{\mathbb{R}}}^{d}\setminus\{x\}$ to $\iota_{x}(a)=x+(a-x)/{\|{a}-{x}\|}^{2}$. It is not difficult to see that the image under this inversion of a hyperplane that avoids $x$ is a sphere that passes through $x$. Similarly, the image of the open half-space that does not contain $x$ is the open ball bounded by this sphere. If the hyperplane passes through the points of a $k$-facet, $\Delta\subseteq A$, and separates $x$ from the $k$-set on the other side, then the $d$-simplex spanned by $x$ and the points in $\iota_{x}(\Delta)$ is a $k$-hefty simplex of $A^{\prime}=\iota_{x}(A)\cup\{x\}$ incident to $x$.
Write $F_{k}(A)$ for the number of $k$-facets of $A$. We first reprove the following $2$-dimensional result by Alon and Györi [1] for $k$ smaller than a third of the number of points.
Proposition 3.3.
Let $A$ be a generic set of $n$ points in ${{\mathbb{R}}}^{2}$ and $k<\frac{n}{3}$ a non-negative integer. Then $\sum_{j=0}^{k}F_{j}(A)\leq(k+1)n$.
Proof 3.4.
Recall that the $k$-hull of $A$ is the intersection of all closed half-spaces that miss at most $k$ points of $A$, denoted ${{H}_{k}{({A})}}$. By the Centerpoint Theorem of discrete geometry, ${{H}_{k}{({A})}}$ has a non-empty interior if $k<\frac{n}{3}$; see e.g. [5, Section 4.1]. Let $x$ be a point in the interior of ${{H}_{k}{({A})}}$ and set $A^{\prime}=\iota_{x}(A)\cup\{x\}$. Let $j\leq k$. As explained above, the inversion through the unit circle centered at $x$ maps every $j$-facet of $A$ to a $j$-hefty triangle of $A^{\prime}$ incident to $x$. By Corollary 2.9, the $j$-hefty triangles incident to $x$ cover a small neighborhood of $x$ at most $j+1$ times, so if we consider all $j$ between 00 and $k$, we get the neighborhood of $x$ covered at most $1+2+\ldots+(k+1)=\binom{k+2}{2}$ times.
To continue, we draw a half-line emanating from $x$ through every point in $\iota_{x}(A)$, thus splitting the neighborhood of $x$ into $n$ angles. Every hefty triangle incident to $x$ covers a contiguous sequence of these angles, and for each $i\geq 1$, there are at most $n$ triangles that cover exactly $i$ of these angles. Each angle is covered some integer number of times, and we take the sum of these numbers over all angles. If $\sum_{j=0}^{k}F_{j}(A)>(k+1)n$, then this sum is strictly greater than $n(1+2+\ldots+(k+1))=n\binom{k+2}{2}$. This implies that one of the angles is covered more than $\binom{k+2}{2}$ times, which contradicts Corollary 2.9.
We continue with Lovász Lemma—see [3] but also [8, Lemma 11.3.2]—which is a crucial ingredient in many arguments about $k$-sets; see e.g. [9]. This lemma gives an asymptotic upper bound on the number of $(d-1)$-simplices spanned by the points of $k$-facets a line can intersect. We give a short proof of a more general version by Welzl [12]. To state the lemma, we say a directed line, $L$, enters a $k$-facet, $\Delta\subseteq A$, if $L$ intersects the $(d-1)$-simplex spanned by $\Delta$ at an interior point and moves from the side with $k$ points to the side with $n-d-k$ points as it passes through the $(d-1)$-simplex.
Proposition 3.5.
Proof 3.6.
We may assume that $L$ passes through the convex hull of $A$ and let $x\in L$ be a point reached after passing through ${{\rm conv\,}{A}}$. The inversion through the unit sphere centered at $x$ maps every $k$-facet entered by $L$ to a $k$-hefty simplex incident to $x$ such that $L$ passes through this simplex before it reaches $x$. By Corollary 2.9, there are at most $\binom{d+k-1}{d-1}$ such $k$-hefty simplices.
Since a line can be directed in two different ways, the number of $(d-1)$-simplices spanned by $k$-facets of the set $A$ one orientation of this line or the other can enter is at most $2\binom{d+k-1}{d-1}$.
@@ 4 Concluding Remarks
The results on covering numbers presented in Section 2 extend to other settings of which we mention three:
  • locally finite sets in hyperbolic space;
  • finite sets on the sphere;
  • weighted points in Euclicean space.
A claim about global covering analogous to Theorem 2.3 holds in all three settings, while one about local covering analogous to Theorem 2.7 holds only for the first two. The first setting is most similar to the situation in Euclidean space studied in Section 2: the proofs are almost verbatim the same, with the main difference being a modified radial set since the design illustrated in Figure 2 is not thin Delone in hyperbolic space. For the second setting, there is an ambiguity in the definition of a $k$-hefty simplex since any $(d-1)$-sphere on the $d$-sphere bounds two complementary $d$-balls. We thus require that the finite set is $k$-balanced, by which we mean that every open hemisphere contains at least $k+1$ points. With this assumption, only the smaller of the two open $d$-balls has a chance to contain $k$ of the points, so we consider it the inside of the $(d-1)$-sphere. The proofs of the extensions of Theorems 2.3 and 2.7 to $k$-balanced sets on the sphere are similar to those for thin Delone sets in Euclidean space. Alternatively, we may use stereographic projection to reduce the spherical to the Euclidean setting.
The generalization to points with real weights is aking to the generalization of Voronoi to power tessellations or diagrams; see e.g. [5, Chapter 13]. We interpret a point with weight $w\in{{\mathbb{R}}}$ as a sphere with squared radius $w$ (which may be negative), and we generalize the circumsphere of a $d$-simplex to the orthosphere, which in the generic case is the unique sphere orthogonal to the $d+1$ spheres that are the vertices of the $d$-simplex. The orthosphere encloses a weighted point if the two centers are strictly closer than required for the two spheres to be orthogonal to each other, and the $d$-simplex is $k$-hefty if it encloses exactly $k$ of the weighted points. With these adaptations of the definitions, a claim analogous to Theorem 2.3 holds also in the weighted case, but Theorem 2.7 fails to extend. Indeed, the $k$-hefty simplices incident to a weighted point cover the neighborhood of the point some integer number of times, but depending on the point and its surrounding, this integer varies between 00 and $\binom{d+k-1}{d-1}$.

### v3

@@ 4.4 Worpitzky’s Identity for Eulerian Numbers
As the fourth and last application of our covering results, we show how the multiplicities from Theorem 2.3 are related to the volumes of hypersimplices and to Eulerian numbers. Specifically, we show that de Laplace’s relation for hypersimplices [14] implies Worpitzky’s identity for Eulerian numbers [25], and vice versa.
We begin by introducing the three main concepts we need in this subsection. Letting $d$ be a positive integer, the number of descents in a permutation $j_{1},j_{2},\ldots,j_{d}$ of $1,2,\ldots,d$ is the number of indices $1\leq i\leq d-1$ such that $j_{i}>j_{i+1}$. For $0\leq k\leq d-1$, the Eulerian number for $d$ and $k$ is the number of permutations of $1,2,\ldots,d$ with exactly $k$ descents. For example, A(d,0)=A(d,d1)=1A(d,0)=A(d,d-1)=1, for every $d$, and $\sum_{k=0}^{d-1}A(d,k)=d!$. Less obvious is Worpitzky’s identity [25] between two polynomials from more than a century ago:
@@
Next, let $x_{0},x_{1},\ldots,x_{d}$ be $d+1$ affinely independent points and write $\Delta={{\rm conv\,}{\{x_{0},x_{1},\ldots,x_{d}\}}}$ for the $d$-simplex they span. A $p$-fold sum is obtained by selecting $p$ of the $d+1$ points, considering them as vectors, and returning the point that corresponds to the sum of these vectors. For each $1\leq p\leq d$, the convex hull of the $p$-fold sums of the $d+1$ points is a $d$-dimensional convex polytope referred to as a $d$-hypersimplex of order $p$, denoted $\Delta_{d}^{(p)}$. Clearly, $\Delta_{d}^{(1)}=\Delta$, and more generally, $\Delta_{d}^{(p)}$ is a homothetic copy of the convex hull of the barycenters of all (p1)(p-1)-dimensional faces of $\Delta$. Since the barycenter of a (p1)(p-1)-simplex is $\nicefrac{{1}}{{p}}$ times the sum of its vertices, the volume of that polytope is $\nicefrac{{1}}{{p^{d}}}$ times the volume of the homothetic hypersimplex. Define the relative volume of $\Delta_{d}^{(p)}$ as $v(d,p)={\operatorname{vol}_{d}(\Delta^{(p)}_{d})}/{\operatorname{vol}_{d}(\Delta)}$, and observe that it does not depend on the choice of $\Delta$. Again more than a century ago, de Laplace [14] proved that these relative volumes are Eulerian numbers:
@@
see also the combinatorial proof of the same equation by Stanley [21] and an interpretation of de Laplace’s result by Foata [12]. The third concept is the dual of the order-$n$ Voronoi tessellation of a finite set $A\subseteq{{\mathbb{R}}}^{d}$, introduced in 1990 by Aurenhammer [5]. Referring to this dual as the order-$n$ Delaunay mosaic of $A$, denoted ${\rm Del}_{n}{({A})}$}, it is defined by its $d$-cells, each the convex hull of a collection of averages of $n$ points selected from $A$. Specifically, for each $1\leq p\leq d$ and every (np)(n-p)-hefty simplex, take all sets of cardinality $n$ that contain all npn-p points inside the circumscribed sphere together with any $p$ points on the circumscribed sphere. E.g. for $p=1$, we get $d+1$ averages whose convex hull is a homothetic copy of the original $d$-simplex, and for $p=2$, we get a homothetic copy of the convex hull of the midpoints of the edges of the $d$-simplex. Collecting these polytopes, we get ${\rm Del}_{n}{({A})}$}. For $n=1$, we have ${{\rm Del}_{1}{({A})}}={{\rm Del}{({A})}}$, and more generally ${\rm Del}_{n}{({A})}$} has a $d$-cell for every (np)(n-p)-hefty simplex in which $p$ varies from $1$ to $d$. Since the vertices are averages of $n$ points, each $d$-cell in ${\rm Del}_{n}{({A})}$} has volume $\nicefrac{{1}}{{n^{d}}}$ times the volume of the corresponding hypersimplex. It is now easy to prove the following relation for the relative volumes of the hypersimplices.
@@
in which $\binom{n+d-p}{n-p}=0$ whenever np<0n-p<0.
Proof 4.11.
Let $A$ be any Delone set—and not just a thin Delone set—in ${{\mathbb{R}}}^{d}$, so that every ball whose radius exceeds some given constant contains at least one point of $A$. Let $R>0$ be sufficiently large and consider all $d$-cells in ${\rm Del}_{n}{({A})}$} that are contained in $R,R]^{d}$. Setting $n^{\prime}=\max\{0,n-d\}$, there are $n-n^{\prime}\leq d$ different types of $d$-cells to be considered, namely homothetic copies of hypersimplices of orders $1$ to $n-n^{\prime}$ defined by (np)(n-p)-hefty simplices for $1\leq p\leq n-n^{\prime}$. The total volume of these $d$-cells is $(2R)^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$, since we miss only a constant width neighborhood of each facet of the hypercube.
Consider now an (np)(n-p)-hefty simplex of $A$. Generically, its circumscribed sphere passes through $d+1$ points and encloses npn-p of the points in $A$. By definition of Delone set, the radius of this circumscribed sphere is bounded from above by a constant times npn-p. Furthermore, the (np)(n-p)-hefty simplex contains every $d$-cell in ${\rm Del}_{n}{({A})}$} it may determine since the vertices of the latter are averages of the vertices of the (np)(n-p)-hefty simplex. By Theorem 2.3, the (np)(n-p)-hefty simplices that define $d$-cells of ${\rm Del}_{n}{({A})}$} inside the hypercube therefore cover most of the hypercube exactly $\binom{n+d-p}{n-p}$ times. It follows that the total volume of these $p$-hefty simplices is $\binom{n+d-p}{n-p}(2R)^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$. By definition of relative volume, the total volume of the corresponding hypersimplices thus is ${v(d,p)}\binom{n+d-p}{n-p}(2R)^{d}/n^{d}+{\mathcal{O}\left({R^{d-1}}\right)}$. Taking the sum for $1\leq p\leq n-n^{\prime}$, dividing by $(2R)^{d}$, and taking the limit as $R$ goes to infinity, we get the claimed relation.
To see that Theorem 4.10 together with de Laplace’s relation implies Worpitzky’s identity and together with Worpitzky’s identity implies de Laplace’s relation, we change the summation index in (7) from $p$ to k=dpk=d-p and apply $\binom{n+k}{n-d+k}=\binom{n+k}{d}$ to get
@@
Since this relation holds for every positive integer, $n$, it also holds if we treat $\binom{n+k}{d}$ and $n^{d}$ as polynomials of degree $d$ in $n$. Substituting A(d,k)=A(d,dk1)A(d,k)=A(d,d-k-1) for v(d,dk)v(d,d-k) using de Laplace’s relation (6), we get Worpitzky’s identity (5). To see the other direction, we observe that the polynomials given by the binomial coefficients are linearly independent, so there is only one way to write $n^{d}$ as their linear combination, and it is given by Worpitzky’s identity. Comparing (8) with (5), we get v(d,dk)=A(d,k)=A(d,dk1)v(d,d-k)=A(d,k)=A(d,d-k-1), which is (6).
@@ 5 Concluding Remarks
Given a finite or locally finite set of points, this paper generalizes Delaunay’s concept of simplices with empty circumscribed spheres [8] to what we call $k$-hefty simplices, and proves that they cover space a predictable number of times. We use this insight to give new proofs of old and new results in discrete geometry, thus providing evidence for the potential of these simplices. Besides points and weighted points in Euclidean space, we study the spherical setting but omit others, like hyperboloic space, to which our results extend as well. In conclusion, we highlight additional connections to topics in computational geometry and combinatorics, and formulate questions we expect can be tackled with the methods studied in this paper.
Maximal Feasible Subsystems. Let $A{\bf{x}}\leq{\bf b}$ be a system of $n$ linear inequalities in ${{\mathbb{R}}}^{d}$ specifying the negative halfspaces considered in Section 4.3. When $k=0$, the system is consistent and the solution set is a convex polyhedron. For $k>0$, the system is inconsistent and we need to remove at least $k$ inequalities to make it consistent.
For a given inconsistent system of inequalities, the problem of finding any minimal subset whose removal makes the system consistent is the Maximal Feasible Sybsystem (Max FS) problem, which is known to be NP-hard [2]. Theorem 4.7 implies that if a generic linear system can be made consistent by the removal of $k$ inequalities, but not by the removal of fewer than $k$ inequalities, then there are at most $\binom{d+k}{d}$ options to do that.
Chambers in Oriented Matroids. Every arrangement of oriented hemispheres gives rise to a realizable oriented matroid with chambers marked by sign vectors, where the $i$-th sign records whether the chamber lies inside or outside the $i$-th hemisphere; see e.g. [18]. Assuming each chamber of a uniform realizable oriented matroid of rank $d+1$ has at least $k$ plus signs, Remark 4.9 gives an upper bound on the number of chambers with exactly $k$ plus signs. We note that this bound is different when the number of hemispheres is arbitrary and when this number is sufficiently large, in which case the bound does not depend on it.
We conjecture that the same bounds hold for the case of uniform non-realizable oriented matroids, or, equivalently, for the case of simple arrangements of pseudospheres in ${{\mathbb{S}}}^{d}$.
More on $k$-facets and $k$-sets. As described in Section 4.1, our theorems on $k$-hefty simplices can be used to re-prove classic bounds on $k$-sets and $k$-facets. However, the authors of this paper have yet been unable to re-prove the tight lower bound on the number of “at most $k$-sets”. Motivated by the connection to the rectilinear crossing number of the complete graph, $K_{n}$, [15] proves $\sum_{j=0}^{k}F_{j}(A)\geq 3\binom{k+2}{2}$ for every generic set $A$ of $n\geq 2k$ points in ${{\mathbb{R}}}^{2}$. For $n\geq 3k+3$, we may assume that $0\in{{\mathbb{R}}}^{2}$ is an interior point of the $k$-hull, so we can invert $A$ through the unit circle centered at 0 and for each $j$-facet get a $j$-hefty triangle incident to 0. By Theorem 2.5, the “at most $k$-hefty” triangles incident to 0 cover any sufficiently small neighborhood of 0 exactly $\binom{k+2}{2}$ times. Since every triangle has an angle strictly less than $\pi$ at 0, we therefore have at least $2\binom{k+2}{2}+1$ such triangles. But the bound in [15] says there are at least $3\binom{k+2}{2}$ such triangles.
The above argument generalizes to proving that there are at least $2\binom{d+k}{d}+1$ “at most $k$-hefty” simplices incident to any one point in a generic set in ${{\mathbb{R}}}^{d}$. This lower bound appears already in [22, Theorem 6.3], along with the conjecture that it can be strengthened to at least $(d+1)\binom{d+k}{k}$ “at most $k$-hefty” simplices in ${{\mathbb{R}}}^{d}$.
From $k$-hefty Simplices to Manifolds. A yet open question motivated by the work in this paper is the global connectivity of the $k$-hefty simplices of a given set. For example, the $k$-hefty triangles of a finite $k$-balanced set in ${{\mathbb{S}}}^{2}$ cover $\binom{k+2}{2}$ times, but this does not necessarily mean that they form that many spheres each covering ${{\mathbb{S}}}^{2}$ once.
By duplicating vertices and edges as needed, the complex of $k$-hefty triangles and their faces can be turned into an orientable $2$-manifold. However, this $2$-manifold is not necessarily a universal cover of ${{\mathbb{S}}}^{2}$ as there may be vertex stars that cover the local neighborhood multiple times. It would be interesting to characterize the manifolds that can arise—in $2$ but also higher dimensions—how they cover, and which local combinatorial properties determine the global connectivity of the manifold.
Acknowledgments
The authors would like to thank Boris Aronov and János Pach for pointing us to result of Clarkson that we reprove in Section 4.2. We also would like to thank Jesús De Loera for a discussion on Maximal Feasible Subsystems.